OpenStudy (waheguru):

How do I convert y = 2/3 x - 7 to standard ax + by +c = 0 form

OpenStudy (waheguru):

*more so I do i get the = 0 part...??

OpenStudy (mathstudent55):

Since you have a denominator of 3, first multiply both sides by 3. Then add or subtract the terms on the right side to move them to the left side.

OpenStudy (mathstudent55):

y = 2/3 x - 7 Multiply both sides by 3: (3)y = (3)(2/3 x - 7) What do you get? Remember to distribute the 3 on the right side.

OpenStudy (waheguru):

so 3y = 2x -21

OpenStudy (waheguru):

what do i do from here?

OpenStudy (mathstudent55):

Now since you want everything on the left side, subtract 2x from both sides, and add 21 to both sides.

OpenStudy (waheguru):

can you please show the following steps

OpenStudy (mathstudent55):

3y = 2x - 21 -2x + 21 -2x + 21 -------------------------- -2x + 3y + 21 = 0 If you don't like to have a negative sign at the beginning (with the x term), you can multiply both sides by -1 and get 2x - 3y - 21 = 0

OpenStudy (waheguru):

but i also saw a formula in standard from as 2x-3y = 12. would this be condisered stard from

OpenStudy (waheguru):

*standard form